Hello, I'm trying to help my son with his math homework but it's been 30 years since I had to do math and I can't remember enough to help him.
One example of the type of questions he has would be:
tennis consumes 25 calories a minute
cycling consumes 35 calories a minute
The player plays both sports for a total of 50 minutes consuming 1450 calories. How long did he play tennis?
I've been trying to come up with something like 25x + 35y = 1450 but I can't go from there.
I'd be grateful if you could quickly give me some idea how to do the above.
Thanks
John Mullany

Answer by Alan3354(69443)

tennis consumes 25 calories a minute
cycling consumes 35 calories a minute
The player plays both sports for a total of 50 minutes consuming 1450 calories. How long did he play tennis?
I've been trying to come up with something like 25x + 35y = 1450 but I can't go from there.
I'd be grateful if you could quickly give me some idea how to do the above.
---------------------------
25x + 35y = 1450 is the right approach, but I'll use t and c
25t + 35c = 1450 (same thing)
t + c = 50 (the missing piece)
------------------
5t + 7c = 290 (reduced, smaller numbers)
Substitute t = 50-c
5*(50-c) + 7c = 290
250 - 5c + 7c = 290
2c = 40
c = 20
--------------
Then t = 30